The differential equation corresponding to all circle x^2+y^2-2ay=a^2,is:

Grade:12th pass

1 Answers

Arun
6 years ago
Dear Shivam
 
On differentiating
 
2 x + 2 y *y' - 2a y' = 0
 
a y'= x + y* y'
 
a = (x + y * y')/y'
 
Now
 
Put a in original circle equation.
 
Hence you will get your answer.
 
 
Regards
Arun (askIITians forum expert)